The Advanced Tech Company has a project to design an integrated information database for a major bank. Data for the project are given in Table 1 below. Indirect project costs amount to $300 per day. The company will incur a $150 per day penalty for each day the project lasts beyond day 14. (a) What is the shortest time duration for this project regardless of cost? 

 (b) What is the total cost and time of the minimum-cost schedule?
Answer #1

ANSWER (a)

Project duration is 23 days i.e. shortest time duration for this project regardless of cost

Project

Activity
time

Predessor

ES

EF

LS

LF

Slack
(LS-ES)

A

6

-

0

6

0

6

0

B

4

-

0

4

2

6

2

C

3

A,B

6

9

6

9

0

D

2

B

4

6

7

9

3

E

6

C,D

9

15

9

15

0

F

2

E

15

17

21

23

6

G

4

E

15

19

15

19

0

H

4

G

19

23

19

23

0



Project

23





فه 0 D Q

Critical Path is A-C-E-G-H = 6+3+6+4+4 = 23

ANSWER (b)

Project

Normal time

Crash time

Normal cost

crash cost

Crash cost/pd

Crash by

Crashing cost

A

6

5

1000

1200

200

1

200

B

4

2

800

2000

600

0

0

C

3

2

600

900

300

1

300

D

2

1

1500

2000

500

0

0

E

6

4

900

1200

150

2

300

F

2

1

1300

1400

100

0

0

G

4

4

900

900

0

0

0

H

4

2

500

900

200

2

400

Totals



7500




1200

We can crash only Critical path activities to shorten project duration.

So Max crash time is 6 days for $1200.

So Minimum Project duration = 23-6 = 17 days

So Penalty is payable for 17-14 = 3 days

So minimum cost = Project cost + Crash cost + Indirect cost + penalty

= 7500 + 1200 + 300*17 + 3*150

= 14,250

Critical path has Not changed & remains A-C-E-G-H

Critical path length is now 5+2+4+4+2 = 17

N 2 F 5 2 А E H 4 2 B D G

Activity A A, B ש ס ס ח ו ם 1 Normal Time ( days) Normal cost ($) Crash time ( days) Crash cost ($) Immediate Predecessor (s)


answered by: ANURANJAN SARSAM